Two trains leave the station at the same time, one heading east and the other west. The eastbound train travels
20
miles per hour slower than the westbound train. If the two trains are
570
miles apart after
3
hours, what is the rate of the eastbound train?
Do not do any rounding.

Answer by lwsshak3(11628)

Two trains leave the station at the same time, one heading east and the other west. The eastbound train travels 20 miles per hour slower than the westbound train. If the two trains are 570 miles apart after 3 hours, what is the rate of the eastbound train?
***
let x=speed of eastbound train
x+20=speed of westbound train
x+(x+20)=2x+20=speed at which the 2 trains are traveling apart
Travel time*speed=distance
3*(2x+20)=570
6x+60=570
6x=510
x=85
What is the rate of the eastbound train? 85 mph